# Non-monotonic specific entropy on the transition line near the QCD critical point

### v1

@@ Abstract
We investigate the effect of the QCD critical point on the isentropic trajectories in the QCD phase diagram. We point out that the universality of the critical equation of state and the third law of thermodynamics require the specific entropy (per baryon) along the coexistence (first-order transition) line to be non-monotonic at least on one side of that line. Specifically, a maximum must occur. We show how the location of the maximum relative to the QCD critical point depends on the parameters of the critical equation of state commonly used in the literature. We then examine how the isentropic trajectories followed by adiabatically expanding heavy-ion collision fireballs behave near the critical point. We find that a crucial role is played by the sign of the isochoric temperature derivative of pressure at the critical point; this sign determines on which side of the coexistence curve the specific entropy must be non-monotonic (i.e., has a maximum). We classify different scenarios of the adiabatic expansion that arise depending on the value of the discriminant parameter and the proximity of the trajectory to the critical point.
@@ I Introduction
Ideal hydrodynamics provides the lowest order approximation of the expanding fireball created in heavy-ion collisions [2]. Due to the scale hierarchy between the size of the fireball and the microscopic QCD scale, the dissipation accompanying the expansion can be considered small and the entropy to be approximately conserved. Since the baryon number is also conserved, the entropy per baryon number $\hat{s}\equiv s/n$, which we refer to as specific entropy, is approximately constant, even though both densities of entropy, $s$, and of baryon number, $n$, decrease due to expansion. Therefore, given the equation of state (EOS), i.e., the dependence of thermodynamic quantities, such as pressure on temperature $T$ and baryon chemical potential $\mu$, one can identify the trajectories on the QCD phase diagram as lines of constant $\hat{s}$.
@@
This universal behavior stems from that of the order parameter field, $\phi\propto\pm\left[(T_{\rm c}-T)/T_{\rm c}\right]^{\beta}$, in the conformal theory describing critical phenomena. The exponent $\beta\approx 0.326$ is universal, while the amplitude of the singularity is given by a non-universal coefficient $B_{\phi}$ depending on the specific details of the theory. The ellipsis refers to subleading terms. [^1] We shall see that when the maximum is sufficiently close to the critical point these subleading terms are responsible for creating that maximum. The third law is not even needed to establish the non-monotonicity in this case. Close to the critical point one of the branches is thus necessarily an increasing function of the distance from the critical point. However, at $T=0$ this function must vanish by the third law. Hence, on this branch, a maximum must occur, as illustrated by the green point on Fig. 1.222While Fig. 1 assumes the simplest scenario where the coexistence line extends all the way to $T=0$, it can be generalized to more sophisticated phase diagram scenarios.
@@
In order to study this effect quantitatively we consider the range of parameters (relevant for the realistic QCD EOS) in which the behavior of $\hat{s}$ near the maximum is determined by the competition between the leading term [shown in Eq. (1)] and the subleading singular terms. We employ the critical EOS proposed in Ref. [3] and reveal the universal properties of the specific entropy, such as the non-monotonic structure along the first-order boundary (coexistence line).
@@
We investigate the topography of the specific entropy as a curved surface on a two-dimensional plane of baryon chemical potential $\mu$ and $T$, $\hat{s}(\mu,\,T)$, and demonstrate the generic presence of a ridge-like structure (see the yellow regions of Figs. 5 and 6). The non-monotonic behavior, i.e., the maximum, along the coexistence line illustrated in Fig. 1 corresponds to the cross-section of that ridge.
@@
Such a ridge and the corresponding maximum on the coexistence line emerge either on the HRG side or on the QGP side of the transition. A simple way to see which side is nonmonotonic is to ask which of the coexisting phases has higher specific entropy at the same temperature (see Fig. 1). The answer to this question depends on the sign of $B_{\phi}$ — the coefficient of the leading singularity in Eq. (1). We determine this coefficient in terms of the parameters characterizing the mapping of the thermodynamic singularity of the 3D Ising model to that of the QCD critical point, introduced in Ref. [3]. We also find that the sign determining the non-monotonic side is the same as that of the temperature derivative of pressure at fixed baryon density [see Eqs. (33) and (34) below].
@@
As an example, using the BEST Collaboration EOS with default mapping parameter set choice from Ref. [3], we find that the ridge giving rise to the maximum of the specific entropy on the coexistence line emerges on the HRG side. For this scenario, temperature decreases and chemical potential increases as the isentropic trajectory traverses the coexistence region. Such a scenario is different from the scenario where the system is “reheated” upon traversing the coexistence region by the entropy released in the process of re-confinement (as in, e.g., Ref. [4]), but it has been observed in some models of the QCD phase transition without confinement (see, e.g., Ref. [5]).
@@
The non-monotonic structure can be characterized by two points: the maximum (green point on Fig. 1) and the point where the specific entropy again equals $\hat{s}_{\rm c}$, which we call “critical double” (gray on Fig. 1). Isentropic trajectories can be classified based on the location of the point where they enter the coexistence region (or, if they enter it at all). An interesting class is represented by trajectories that enter the coexistence region between the critical point and its double. Such trajectories emerge on the same side of the coexistence line (see Fig. 8 below). [^3] Such trajectories could also be seen in the $s$ vs $n$ plane in Ref. [6] and possibly on the $(\mu\,,T)$ plane in Refs. [7, 8], although not much attention was given to them.
@@
A series of appendices complements our analysis. Appendix A lists the critical susceptibilities used in Secs. III and V. Appendix B reviews the EOS of the 3D Ising universality class. Appendix C derives specific values of the scaling functions utilizing the findings from Appendix B. Appendix D reviews the mathematical tools used to explore the topography of $\hat{s}$ on a two-dimensional plane, including the ridge line definition. Appendix E analyses the specific heat singularity and relates it to the singularity in the specific entropy expressed in Eq. (1).
@@ II.1 The map
Figure 2: Illustration of the role of the angles $\alpha_{1}$ and $\alpha_{2}$ in the mapping of the Ising model coordinates $(h,\,r)$ onto the phase diagram of QCD in $(\mu,\,T)$ plane given by Eq. (II.1). Panels (b)-(d) show the mapping of the points in the four quadrants (color coded) of the $(h,\,r)$ plane onto $(\mu,\,T)$ plane under two different scenarios for the angle difference $\alpha_{12}=\alpha_{1}-\alpha_{2}$ used in the literature (see text).
@@
We illustrate the mapping geometrically in Fig. 2 to compare and contrast two scenarios different by the angle difference $\alpha_{12}$. The default scenario choice of Ref. [3], often used in the literature, is $\alpha_{12}=90^{\circ}$ in panel (c). There is no motivation for this choice apart from simplicity. On the other hand, as pointed out in Ref. [10], in the massless quark limit, $m_{q}\to 0$, the angle difference $\alpha_{12}$ vanishes. [^5] The mapping becomes singular (non-invertible) in the $\alpha_{12}\to 0$ limit. This reflects the fact that the criticality in this limit corresponds to a tricritical point, in a different universality class [4, 11, 10]. Specifically: $\alpha_{12}\to+0$ as $\alpha_{12}\sim m_{q}^{2/5}$.
This fact, together with the smallness of quark masses, suggests that $\alpha_{12}$ is small and $\alpha_{2}<\alpha_{1}$ in Nature. This is the scenario illustrated in Fig. 2, panel (d).
@@ II.2 Specific entropy
with normalization factor $A=T_{\rm c}^{4}$ as in Ref. [3]. The $\dots$” denotes less singular (and regular) terms also vanishing at the critical point. These terms are negligible, sufficiently close to the critical point compared to the singular terms we write out. Specifically, along the coexistence line, $\phi\sim|r|^{\beta}$ and $\varepsilon\sim|r|^{1-\alpha}$, while the omitted terms are at most of order $|r|$. [^6] Given the values of the exponents $\beta\approx 0.326$ and $\alpha\approx 0.110$, for sufficiently small $|r|$, the hierarchy $|r|^{\beta}\gg|r|^{1-\alpha}\gg|r|$ holds. We shall focus on such a regime and revisit the justification for neglecting the $\propto|r|$ term in Sec. III.5. The leading and subleading singularities in Eq. (II.2) are due to the two relevant operators of the conformal $\phi^{4}$ theory corresponding to the magnetization, or order parameter, $\phi$ and the energy density $\varepsilon$ [12]:
@@
whose $(h,\,r)$ dependence will be discussed further later. Here, the subscripts denote partial derivatives with respect to one of the independent variables in a set, such as $(\mu,\,T)$ or $(h,\,r)$, while the other variable in the set is fixed. We have used the chain rule, $G_{X}=G_{h}h_{X}+G_{r}r_{X}$ to obtain the last expression of Eqs. (4b) and (4c). Partial derivatives of $h$ and $r$ with respect to $\mu$ and $T$ remain constant due to the linearity of the mapping (II.1):
@@
In Eq. (II.2), $P_{\rm c}$, $s_{\rm c}$, and $n_{\rm c}$ represent the constant baseline values at the critical point. They are, of course, non-universal and later, in Sec. III.4, we will be using the values obtained from extrapolating lattice data incorporated in the BEST Collaboration EOS in Ref. [3].
@@
The coefficients $(\hat{s}_{\phi})_{\rm c}$ and $(\hat{s}_{\varepsilon})_{\rm c}$, in particular, their signs, will play a central role in determining the global structure of $\hat{s}$.
@@
${\rm sgn}\,(\hat{s}_{\phi})_{\rm c}$ $={\rm sgn}\,(\hat{s}_{\rm c}-\cot\alpha_{1})\,,$ (9a)
${\rm sgn}\,(\hat{s}_{\varepsilon})_{\rm c}$ $=-{\rm sgn}\,(\hat{s}_{\rm c}-\cot\alpha_{2})\,.$ (9b)
@@ III.1 $(r,\,h)$ plane
Figure 3: Specific entropy along the coexistence line given by Eq. (14). The black curve represents the branch demonstrating a non-monotonic behavior, which is one of the main results in the present paper. The dashed curve is the branch that remains monotonic. The black, gray, and green points denote the critical point, critical double, and the maximum of the non-monotonic branch (see text).
@@
where the coefficient $(\hat{s}_{\phi})_{\rm c}$ vanishes (and changes sign), according to Eq. (II.2). At this point $r_{\rm max}$ and $r_{\rm cd}$ vanish. This is the regime (sufficiently small $|r|$) where our quantitative analysis based on the leading critical behavior of the EOS applies. It should be kept in mind, however, that even outside of this regime, the existence of the maximum is guaranteed, as discussed in the Introduction, by the third law of thermodynamics.
@@
While angle $\alpha_{1}$ controls the magnitude of $r_{\rm max}$ in Eq. (15) via $(\hat{s}_{\phi})_{\rm c}$, angle $\alpha_{2}$ also affects $r_{\rm max}$, via $(\hat{s}_{\varepsilon})_{\rm c}$. Using Eq. (II.2) we find
@@ III.2 $(\mu,\,T)$ plane
We shall now translate the results of the previous section into the $(\mu,\,T)$ plane using the mapping Eq. (II.1). On the coexistence line, $h=0$, Eq. (II.1) reduces to
@@
III.3 The non-monotonic branch
As we discussed before, the branch exhibiting the maximum is the one where $\hat{s}$ is increasing away from the critical point (see Fig. 1). To determine whether it is the high-$T$ (QGP) or the low-$T$ (HRG) branch, we turn back to Eq. (7). The value of $\hat{s}$ is increasing on the branch where $(\hat{s}_{\phi})_{\rm c}\phi>0$. With the choice of the direction of the $h$ axis shown in Fig. 2, the HRG branch corresponds to $h>0$. Thus, given $\phi_{h}=G_{hh}>0$ (thermodynamic stability), $\phi>0$ on this branch. Hence, HRG branch ($h\to+0$) is non-monotonic when $(\hat{s}_{\phi})_{\rm c}>0$. Using Eq. (9a) we conclude:
@@ III.3 The non-monotonic branch
The quantity determining the non-monotonic side of the coexistence line in the $(\mu,\,T)$ plane can be expressed in terms of the isochoric temperature derivative of pressure at the critical point,
@@
This can be seen as a consequence of the fact that the critical isochore ($n=n_{\rm c}$ line for $T>T_{\rm c}$ is tangential to the coexistence line ($h=0$ for $r<0$) at the critical point, i.e., [^8] Indeed, the dominant term in $n-n_{\rm c}$ is proportional to $\phi$, according to Eq. (II.2). Thus the line of constant $n-n_{\rm c}=0$ is the same as the line of constant $\phi=0$, or $h=0$, at the critical point.

### v2

# Nonmonotonic specific entropy on the transition line near the QCD critical point

@@ Abstract
We investigate the effect of the quantum chromodynamics (QCD) critical point on the isentropic trajectories in the QCD phase diagram. We point out that the universality of the critical equation of state and the third law of thermodynamics require the specific entropy (per baryon) along the coexistence (first-order transition) line to be nonmonotonic at least on one side of that line. Specifically, a maximum must occur. We show how the location of the maximum relative to the QCD critical point depends on the parameters of the critical equation of state commonly used in the literature. We then examine how the isentropic trajectories followed by adiabatically expanding heavy-ion collision fireballs behave near the critical point. We find that a crucial role is played by the sign of the isochoric temperature derivative of pressure at the critical point; this sign determines on which side of the coexistence curve the specific entropy must be nonmonotonic (i.e., has a maximum). We classify different scenarios of the adiabatic expansion that arise depending on the value of the discriminant parameter and the proximity of the trajectory to the critical point.
@@ I Introduction
Ideal hydrodynamics provides the lowest-order approximation of the expanding fireball created in heavy-ion collisions [2]. Due to the scale hierarchy between the size of the fireball and the microscopic QCD scale, the dissipation accompanying the expansion can be considered small and the entropy to be approximately conserved. Since the baryon number is also conserved, the entropy per baryon number $\hat{s}\equiv s/n$, which we refer to as specific entropy, is approximately constant, even though both densities of entropy, $s$, and of baryon number, $n$, decrease due to expansion. Therefore, given the equation of state (EOS), i.e., the dependence of thermodynamic quantities, such as pressure on temperature $T$ and baryon chemical potential $\mu$, one can identify the trajectories on the QCD phase diagram as lines of constant $\hat{s}$.
@@
This universal behavior stems from that of the order parameter field, $\phi\propto\pm\left[(T_{\rm c}-T)/T_{\rm c}\right]^{\beta}$, in the conformal theory describing critical phenomena. The exponent $\beta\approx 0.326$ is universal, while the amplitude of the singularity is given by a nonuniversal coefficient $B_{\phi}$ depending on the specific details of the theory. The ellipsis refers to subleading terms. [^1] We shall see that when the maximum is sufficiently close to the critical point these subleading terms are responsible for creating that maximum. The third law is not even needed to establish the nonmonotonicity in this case. Close to the critical point one of the branches is thus necessarily an increasing function of the distance from the critical point. However, at $T=0$ this function must vanish by the third law. Hence, on this branch, a maximum must occur, as illustrated by the green point on Fig. 1.222While Fig. 1 assumes the simplest scenario where the coexistence line extends all the way to $T=0$, it can be generalized to more sophisticated phase diagram scenarios.
@@
In order to study this effect quantitatively we consider the range of parameters (relevant for the realistic QCD EOS) in which the behavior of $\hat{s}$ near the maximum is determined by the competition between the leading term [shown in Eq. (1)] and the subleading singular terms. We employ the critical EOS proposed in Ref. [3] and reveal the universal properties of the specific entropy, such as the nonmonotonic structure along the first-order boundary (coexistence line).
@@
We investigate the topography of the specific entropy as a curved surface on a two-dimensional plane of baryon chemical potential $\mu$ and $T$, $\hat{s}(\mu,\,T)$, and demonstrate the generic presence of a ridge-like structure (see the yellow regions of Figs. 5 and 6). The nonmonotonic behavior, i.e., the maximum, along the coexistence line illustrated in Fig. 1 corresponds to the cross section of that ridge.
@@
Such a ridge and the corresponding maximum on the coexistence line emerge either on the HRG side or on the QGP side of the transition. A simple way to see which side is nonmonotonic is to ask which of the coexisting phases has higher specific entropy at the same temperature (see Fig. 1). The answer to this question depends on the sign of $B_{\phi}$ — the coefficient of the leading singularity in Eq. (1). We determine this coefficient in terms of the parameters characterizing the mapping of the thermodynamic singularity of the three-dimensional (3D) Ising model to that of the QCD critical point, introduced in Ref. [3]. We also find that the sign determining the nonmonotonic side is the same as that of the temperature derivative of pressure at fixed baryon density [see Eqs. (33) and (34) below].
@@
As an example, using the BEST Collaboration EOS with default mapping parameter set choice from Ref. [3], we find that the ridge giving rise to the maximum of the specific entropy on the coexistence line emerges on the HRG side. For this scenario, temperature decreases and chemical potential increases as the isentropic trajectory traverses the coexistence region. Such a scenario is different from the scenario where the system is “reheated” on traversing the coexistence region by the entropy released in the process of reconfinement (as in, e.g., Ref. [4]), but it has been observed in some models of the QCD phase transition without confinement (see, e.g., Ref. [5]).
@@
The nonmonotonic structure can be characterized by two points: the maximum (green point on Fig. 1) and the point where the specific entropy again equals $\hat{s}_{\rm c}$, which we call “critical double” (gray on Fig. 1). Isentropic trajectories can be classified based on the location of the point where they enter the coexistence region (or, if they enter it at all). An interesting class is represented by trajectories that enter the coexistence region between the critical point and its double. Such trajectories emerge on the same side of the coexistence line (see Fig. 8 below). [^3] Such trajectories were seen in the $s$ vs $n$ plane in Ref. [6] (see also Appendix F) and possibly on the $(\mu\,,T)$ plane in Refs. [7, 8], although not much attention was given to them.
@@
A series of appendices complements our analysis. Appendix A lists the critical susceptibilities used in Secs. III and V. Appendix B reviews the EOS of the 3D Ising universality class. Appendix C derives specific values of the scaling functions utilizing the findings from Appendix B. Appendix D reviews the mathematical tools used to explore the topography of $\hat{s}$ on a two-dimensional plane, including the ridge line definition. Appendix E analyses the specific heat singularity and relates it to the singularity in the specific entropy expressed in Eq. (1). In Appendix F, we present a geometric description of the maximum specific entropy phenomenon and the classification of isentropic trajectories in the $(n,\,s)$ plane.
@@ II.1 The map
Figure 2: Illustration of the role of the angles $\alpha_{1}$ and $\alpha_{2}$ in the mapping of the Ising model coordinates $(h,\,r)$ onto the phase diagram of QCD in $(\mu,\,T)$ plane given by Eq. (2). Panels (b)-(d) show the mapping of the points in the four quadrants (color coded) of the $(h,\,r)$ plane onto $(\mu,\,T)$ plane under two different scenarios for the angle difference $\alpha_{12}=\alpha_{1}-\alpha_{2}$ used in the literature (see text).
@@
We illustrate the mapping geometrically in Fig. 2 to compare and contrast two scenarios different by the angle difference $\alpha_{12}$. The default scenario choice of Ref. [3], often used in the literature, is $\alpha_{12}=90^{\circ}$ in Fig. 2(c). There is no motivation for this choice apart from simplicity. On the other hand, as pointed out in Ref. [10], in the massless quark limit, $m_{q}\to 0$, the angle difference $\alpha_{12}$ vanishes. [^5] The mapping becomes singular (noninvertible) in the $\alpha_{12}\to 0$ limit. This reflects the fact that the criticality in this limit corresponds to a tricritical point, in a different universality class [4, 11, 10]. Specifically: $\alpha_{12}\to+0$ as $\alpha_{12}\sim m_{q}^{2/5}$.
This fact, together with the smallness of quark masses, suggests that $\alpha_{12}$ is small and $\alpha_{2}<\alpha_{1}$ in Nature. This is the scenario illustrated in Fig. 2(d).
@@ II.2 Specific entropy
with normalization factor $A=T_{\rm c}^{4}$ as in Ref. [3]. The $\dots$” denotes less singular (and regular) terms also vanishing at the critical point. These terms are negligible, sufficiently close to the critical point compared to the singular terms we write out. Specifically, along the coexistence line, $\phi\sim|r|^{\beta}$ and $\varepsilon\sim|r|^{1-\alpha}$, while the omitted terms are at most of order $|r|$. [^6] Given the values of the exponents $\beta\approx 0.326$ and $\alpha\approx 0.110$, for sufficiently small $|r|$, the hierarchy $|r|^{\beta}\gg|r|^{1-\alpha}\gg|r|$ holds. We shall focus on such a regime and revisit the justification for neglecting the $\propto|r|$ term in Sec. III.5. The leading and subleading singularities in Eq. (4) are due to the two relevant operators of the conformal $\phi^{4}$ theory corresponding to the magnetization, or order parameter, $\phi$ and the energy density $\varepsilon$ [12]:
@@
whose $(h,\,r)$ dependence will be discussed further later. Here, the subscripts denote partial derivatives with respect to one of the independent variables in a set, such as $(\mu,\,T)$ or $(h,\,r)$, while the other variable in the set is fixed. We have used the chain rule, $G_{X}=G_{h}h_{X}+G_{r}r_{X}$ to obtain the last expression of Eqs. (4b) and (4c). Partial derivatives of $h$ and $r$ with respect to $\mu$ and $T$ remain constant due to the linearity of the mapping (2):
@@
In Eq. (4), $P_{\rm c}$, $s_{\rm c}$, and $n_{\rm c}$ represent the constant baseline values at the critical point. They are, of course, nonuniversal and later, in Sec. III.4, we will be using the values obtained from extrapolating lattice data incorporated in the BEST Collaboration EOS in Ref. [3].
@@
The coefficients $(\hat{s}_{\phi})_{\rm c}$ and $(\hat{s}_{\varepsilon})_{\rm c}$, and in particular their signs, will play a central role in determining the global structure of $\hat{s}$.
@@
${\rm sgn}\,(\hat{s}_{\phi})_{\rm c}$ $={\rm sgn}\,(\hat{s}_{\rm c}-\cot\alpha_{1})\cdot\,{\rm sgn}\,(\sin\alpha_{1})\,,$ (9a)
${\rm sgn}\,(\hat{s}_{\varepsilon})_{\rm c}$ $=-{\rm sgn}\,(\hat{s}_{\rm c}-\cot\alpha_{2})\cdot\,{\rm sgn}\,(\sin\alpha_{2})\,.$ (9b)
@@ III.1 $(r,\,h)$ plane
Figure 3: Specific entropy along the coexistence line given by Eq. (14). The black curve represents the branch demonstrating a nonmonotonic behavior, which is one of the main results in the present paper. The dashed curve is the branch that remains monotonic. The black, gray, and green points denote the critical point, critical double, and the maximum of the nonmonotonic branch (see text).
@@
where the coefficient $(\hat{s}_{\phi})_{\rm c}$ vanishes (and changes sign), according to Eq. (8). At this point $r_{\rm max}$ and $r_{\rm cd}$ vanish. This is the regime (sufficiently small $|r|$) where our quantitative analysis based on the leading critical behavior of the EOS applies. It should be kept in mind, however, that even outside of this regime, i.e., when ($|r_{\rm max}|$ is not small), the existence of the maximum is guaranteed, as discussed in the Introduction, by the third law of thermodynamics.
@@
While angle $\alpha_{1}$ controls the magnitude of $r_{\rm max}$ in Eq. (15) via $(\hat{s}_{\phi})_{\rm c}$, angle $\alpha_{2}$ also affects $r_{\rm max}$, via $(\hat{s}_{\varepsilon})_{\rm c}$. Using Eq. (8) we find
@@ III.2 $(\mu,\,T)$ plane
We shall now translate the results of the previous section into the $(\mu,\,T)$ plane using the mapping Eq. (2). On the coexistence line, $h=0$, Eq. (2) reduces to
@@
III.3 The nonmonotonic branch
As we discussed before, the branch exhibiting the maximum is the one where $\hat{s}$ is increasing away from the critical point (see Fig. 1). To determine whether it is the high-$T$ (QGP) or the low-$T$ (HRG) branch, we turn back to Eq. (7). The value of $\hat{s}$ is increasing on the branch where $(\hat{s}_{\phi})_{\rm c}\phi>0$. With the choice of the direction of the $h$ axis shown in Fig. 2, the HRG branch corresponds to $h>0$. Thus, given $\phi_{h}=G_{hh}>0$ (thermodynamic stability), $\phi>0$ on this branch. Hence, HRG branch ($h\to+0$) is nonmonotonic when $(\hat{s}_{\phi})_{\rm c}>0$. Using Eq. (9a) we conclude:
@@ III.3 The nonmonotonic branch
The quantity determining the nonmonotonic side of the coexistence line in the $(\mu,\,T)$ plane can be expressed in terms of the isochoric temperature derivative of pressure at the critical point,
@@
This can be seen as a consequence of the fact that the critical isochore ($n=n_{\rm c}$ line for $T>T_{\rm c}$ is tangential to the coexistence line ($h=0$ for $r<0$) at the critical point, i.e., [^8] Indeed, the dominant term in $n-n_{\rm c}$ is proportional to $\phi$, according to Eq. (4). Thus the line of constant $n-n_{\rm c}=0$ is the same as the line of constant $\phi=0$, or $h=0$, at the critical point.